An exam worth 145 points contains 50 questions. Some of the questions are worth two points and some are worth five points. How many two-point questions are on the test? How many five-point questions are on the test? Write a linear system of equations that can be used to solve this problem. Please help. Thank you!
Answer by Boreal(15235)

x are two point and 50-x are 5 point questions
2x+5(50-x)=145; point value* number of questions=145
2x+250-5x=145
-3x=-105
x=35 2 point questions or 70 points
50-x=15 5 point questions or 75 points. Total is sum or 145 points.